OpenStudy (anonymous):

Could you help me with finding the sum of 7+77+777+...+777...77? Where in the 777...77 are n sevens.

OpenStudy (asnaseer):

The link that @anusha.p pasted above seems to have the solution to this.

OpenStudy (anonymous):

7[10^(n + 1) - 9n - 10]/81. substitute n=7

OpenStudy (anonymous):

Thanks everyone:)

OpenStudy (aravindg):

the question itself cannot be used to find sum , instead we need to make it to form whose sum which we can calculate .I call this process beautification, so here by "beautifying" the expression: \[\large 7+77+777+777+7777+.......\] \[\large =\frac{7}{9}\times(9+99+999+9999+.....n terms)\] \[\large =\frac{7}{9}\times((10^1-1+10^2-1+10^3-1+......2n \;terms)\] \[\large =\frac{7}{9}\times((10^1+10^2+10^3+.....+10^n)-(1+1+1+1+...+n\;terrms))\] \[\large =\frac{7}{9}\times\frac{10(10^n-1)}{10-1})-(n)\] make further simplification if u need :) srry i am slow at typing

OpenStudy (aravindg):

@viodhora?

OpenStudy (anonymous):

thanks a lot:)